### v3

# Factorized Motion Fields for Fast Sparse Input Dynamic View Synthesis

Abstract.
Designing a 3D representation of a dynamic scene for fast optimization and rendering is a challenging task. While recent explicit representations enable fast learning and rendering of dynamic radiance fields, they require a dense set of input viewpoints. In this work, we focus on learning a fast representation for dynamic radiance fields with sparse input viewpoints. However, the optimization with sparse input is under-constrained and necessitates the use of motion priors to constrain the learning. Existing fast dynamic scene models do not explicitly model the motion, making them difficult to be constrained with motion priors. We design an explicit motion model as a factorized 4D representation that is fast and can exploit the spatio-temporal correlation of the motion field. We then introduce reliable flow priors including a combination of sparse flow priors across cameras and dense flow priors within cameras to regularize our motion model. Our model is fast, compact and achieves very good performance on popular multi-view dynamic scene datasets with sparse input viewpoints. The source code for our model can be found on our project page: https://nagabhushansn95.github.io/publications/2024/RF-DeRF.html.
Fast dynamic view synthesis, dynamic radiance fields, factorized models, motion priors, sparse input views
submissionid: 969journal: TOGjournalyear: 2024copyright: rightsretainedconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24; July 27-August 1, 2024; Denver, CO, USAbooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24 (SIGGRAPH Conference Papers ’24), July 27-August 1, 2024, Denver, CO, USAdoi: 10.1145/3641519.3657498isbn: 979-8-4007-0525-0/24/07ccs: Computing methodologies Renderingccs: Computing methodologies Volumetric modelsccs: Computing methodologies Computer visionccs: Computing methodologies Computational photographyccs: Computing methodologies Regularization
1. Introduction
Synthesizing novel views of a scene given images/videos of the scene in other viewpoints has several applications such as meta-verse, sports, and remote exploration. Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) brought in a seminal shift in novel view synthesis by incorporating differential volume rendering with a compact continuous-depth model. However, it has several limitations such as the inability to handle object motion in the scene as well as long training and rendering times. Recently, K-Planes (Fridovich-Keil et al., 2023) significantly reduced the time complexity for dynamic view synthesis by proposing a factorized 4D volume representation. However, K-Planes requires a large number of input viewpoints to render photo-realistic novel views when employing a multi-view camera setup. The goal of our work is to overcome this limitation and design a fast dynamic radiance field that can effectively learn the dynamic scene with few input viewpoints. We mainly focus on a sparse multi view camera setting, where a video from each viewpoint is available.
Optimizing a dynamic radiance field with few input viewpoints is highly under-constrained. A popular approach to regularize an under-constrained system is to impose additional priors during the optimization. For example, sparse static radiance fields are regularized with depth (Deng et al., 2022; Uy et al., 2023), visibility (Somraj and Soundararajan, 2023) and natural image (Wu et al., 2024) priors. However, the object motion in dynamic scenes motivates the study of motion priors for sparse input dynamic view synthesis. Since K-Planes employs a 4D volumetric representation without a motion model, it does not allow the motion implicitly learned by the model to be regularized using motion priors. Thus, there is a need to design a dynamic radiance field with an explicit motion field that lends itself to be constrained with motion supervision.
We design a dynamic radiance field consisting of two models, a 5D radiance field that learns the 3D scene at a canonical time instant and a 4D motion or deformation field that learns the motion from any time instant $t$ to the canonical time instant $t^{\prime}$. Since the motion model is a one-directional mapping from $t$ to $t^{\prime}$, it is not obvious how to impose flow priors across two arbitrary time instants. We achieve this by constraining our motion model to map a pair of matched points, obtained using the motion prior, to the same 3D point in the canonical volume. Fig. 2 illustrates our approach of imposing the motion prior. This allows us to impose motion priors across any two time instants and across any cameras.
Prior and concurrent works (Fang et al., 2022; Shaw et al., 2023) employ deep neural networks (DNN) to learn the motion field. However, the use of the DNN to learn the motion makes the model computationally expensive. The challenge here is to design a motion model that can learn the motion efficiently while yielding fast training and rendering. While explicit models such as voxel grids (Fridovich-Keil et al., 2022; Sun et al., 2022), hash-grids (Müller et al., 2022) and 3D Gaussians (3DGS) (Kerbl et al., 2023) are shown to be effective in learning static scenes, naively extending these techniques to model 4D motion may not be efficient. For example, extending 3D voxel grids to 4D scales the memory requirement to the fourth power of grid resolution. Further, for a given object, since the motion exists at every time instant, models that exploit scene sparsity to reduce the memory requirement may not extend effectively to 4D. Since the motion is dense and has a high spatio-temporal correlation, we employ factorized volumes to exploit the correlation. Specifically, we employ a 4D factorized volume to learn the scene flow from any time $t$ to the canonical time $t^{\prime}$.
An ideal motion prior to regularize the motion field is a dense flow prior across cameras and time instants. Prior models such as NSFF (Li et al., 2021) use dense optical flow to supervise an auxiliary task of predicting the flow between two frames, which may not efficiently regularize the core dynamic radiance field model. Further, we observe that deep flow estimation networks used in such works are sensitive to variations in camera parameters, lighting changes and occlusions, and hence find it challenging to determine reliable dense correspondences across cameras. Regularizing the model with such noisy flow priors may lead to sub-optimal performance.
We address these challenges by employing a combination of two complementary flow priors, a rich but sparse flow prior across cameras and a dense flow prior within individual cameras. In particular, we obtain reliable flow priors across cameras and time instants for a sparse set of SIFT keypoints. Within individual cameras, we obtain reliable dense flow priors using deep optical flow estimation networks (Teed and Deng, 2020). These priors within the same camera tend to be more reliable owing to smaller variations in camera parameters and lighting changes. The reliability of our priors as opposed to the dense flow priors across cameras is illustrated in Fig. 3. The combination of the sparse across camera flow priors and the dense within camera flow priors provides an effective alternative to the cross camera dense flow priors.
We evaluate our model on two popular multi-view dynamic scene datasets and find that our model outperforms the state-of-the-art dynamic view synthesis models with fewer input viewpoints. We refer to our model as RF-DeRF, since we employ Reliable Flow priors for Deformable Radiance Fields. We summarize the main contributions of our work in the following:
  • We design a fast and compact dynamic radiance field for sparse input dynamic view synthesis by employing an explicit motion model that can be easily regularized using motion priors. We employ a 4D factorized volume to exploit the spatio-temporal correlation of the motion field.
  • We propose a complimentary set of reliable flow priors to regularize the motion field when training with few multi-view videos. We obtain a sparse flow prior that regularizes flow across cameras and a dense flow prior that regularizes the flow within individual cameras.
  • We achieve very good dynamic view synthesis performance on two popular multi-view datasets with very few views.
Dynamic Scenes Sparse Views Explicit Model Motion Model
NeRF
TensoRF, i-ngp, 3DGS
DS-NeRF, SimpleNeRF
NSFF, DyNeRF
D-NeRF
TiNeuVox, SWAGS
K-Planes, HexPlane
Ours
Table 1. Related work overview: We compare our work with prior works based on various aspects. Sparse input views refers to models that can handle data from few stationary multi-view cameras. Explicit refers to models that primarily employ explicit models followed by an optional tiny MLP.
TiNeuVox and SWAGS use an explicit model to represent the scene, but use implicit neural networks to model the motion. NSFF predicts motion as an auxiliary task.
2. Related Work
2.1. Classical Work on Deformation Models
The modeling of dynamic radiance fields through a static radiance field and a motion field is similar to the use of a deformation model that deforms a canonical representation of an entity. This is a popular approach to model deformable solids (Sederberg and Parry, 1986), human motion (Loper et al., 2015), facial expressions (Bradley et al., 2010), deformable garments (Miguel et al., 2012; Pons-Moll et al., 2017), fluids (Adrian, 1991; Gregson et al., 2012; Hawkins et al., 2005), gases (Atcheson et al., 2008), smoke (Hawkins et al., 2005), flames (Ihrke and Magnor, 2004) as well as CT (Computed Tomography) and MRI (Magnetic Resonance Imaging) scans in the medical field (Nunes et al., 2012). Learning the motion or deformation field is also found to improve the reconstruction of 3D scenes (Vedula et al., 2000) and free-viewpoint rendering of dynamic scenes (Carranza et al., 2003).
The canonical space in such deformable models is popularly represented as triangular meshes estimated through a multi-view stereo algorithm (Bradley et al., 2008, 2010; Kraevoy and Sheffer, 2005; Allen et al., 2003), or specialized parametric meshes for humans and faces (Anguelov et al., 2005; Vlasic et al., 2008; Loper et al., 2015), or sums of Gaussians (Stoll et al., 2011). In contrast, we employ a radiance field to learn the canonical space owing to its advantages over meshes (Mildenhall et al., 2020). Different deformation models include free-form deformation (FFD) fields or optical flows (Faloutsos et al., 1997; Wang et al., 2009), bijective mappings (Kraevoy and Sheffer, 2004), morphable models (Blanz and Vetter, 1999), splines (Vlasic et al., 2009), volumetric Laplacian deformations (de Aguiar et al., 2008) and linear systems with basis functions (Ihrke and Magnor, 2004). Our motion field is closer to the free-form deformation among the above deformation approaches, but differs in the representation used for the motion field. Finally, the sparse flow priors we employ can be thought of as similar to tracking the markers (Guskov et al., 2003; Hasler et al., 2006) or keypoints (Bradley et al., 2008; Joo et al., 2018) while estimating the classical deformation models.
2.2. Dynamic View Synthesis
Different from the use of deformable models, Zitnick et al. (2004) learn the 3D scene dynamics using a layered depth representation with motion compensation. More recent volumetric representations such as Multiplane Images (MPI) (Zhou et al., 2018) are extended to handle dynamic scenes by employing an MPI per frame (Lin et al., 2021) or temporal basis functions (Xing and Chen, 2021). However, such approaches suffer from depth discretization artifacts (Mildenhall et al., 2020). Instead of volumetric models, Yoon et al. (2020) employ depth image based warping, where the depth is obtained by combining monodepth with multi-view depth. Recent work on dynamic view synthesis with sparse input viewpoints require depth information obtained through the use of RGB-D cameras (Li et al., 2023) or multi-view stereo (Bansal and Zollhöfer, 2023). Further, such approaches struggle to handle soft edges and translucent objects (Penner and Zhang, 2017).
2.3. Dynamic Radiance Fields
In contrast to the classical approaches, dynamic view synthesis can be solved by learning a 6D radiance field that maps the position, time and viewing direction to the radiance. However, the performance of such approaches degrades when the input viewpoints are sparse. We provide a quick overview of prior work that share a few attributes of our model in Tab. 1. To the best of our knowledge, ours is the first work to address dynamic radiance fields with few input viewpoints while achieving fast training and rendering. In the following, we review the literature on dynamic radiance fields and sparse-input static radiance fields.
Dynamic radiance fields can be broadly classified into two categories based on how the temporal modeling is handled, which is crucial in the sparse input setting. A simple approach is to model the dynamic radiance field as a 6D function of position, time, and viewing direction. To utilize flow priors, NSFF (Li et al., 2021) predicts flow as an auxiliary task and constrains the volume density of the mapped 3D points provided by the flow. K-Planes (Fridovich-Keil et al., 2023) and HexPlane (Cao and Johnson, 2023) extend TensoRF (Chen et al., 2022) to a 4D model that maps the position and time to a latent feature, which is then decoded by a tiny multi-layer perceptron (MLP). Instead of conditioning the radiance field directly on time, DyNeRF (Li et al., 2022) conditions the radiance field on a per-time-instant feature vector which is jointly optimized with the radiance field. The lack of a motion model in these approaches makes it incompatible to impose motion priors when learning with sparse input viewpoints.
The second set of models employs a motion or deformation field that maps the 3D points from a given time instant to a canonical time instant (Pumarola et al., 2021; Wang et al., 2023b, a). TiNeuVox (Fang et al., 2022), SWAGS (Shaw et al., 2023) and CoGS (Yu et al., 2023) replace the scene representation MLP in D-NeRF (Pumarola et al., 2021) with a TensoRF or 3DGS model, but use MLPs to model the motion field, leading to expensive training and rendering time. To achieve fast optimization and rendering, DeVRF (Liu et al., 2022) employs a 4D voxel grid to learn the deformation field, but its memory requirements scale with the fourth power of the grid resolution. Instead of employing a free-form deformation model, Wang et al. (2021) models the motion using discrete cosine transform (DCT) basis functions. Different from the above, prior (Guo et al., 2023) and concurrent work (Wu et al., 2023) employ a motion model that maps the 3D points from the canonical time instant to a given time instant. However, the forward warping of points can lead to holes in novel views, which may need to be infilled using a separate inpainting network (Guo et al., 2023; Mirzaei et al., 2023; Somraj et al., 2022).
2.4. Radiance Fields with Sparse Views
Optimizing a radiance field with sparse input views for static scenes has rich literature. With sparse inputs, the volume rendering equations are severely under-constrained leading to distortions such as floaters. Few-shot NeRF models address this by constraining the model further with hand-crafted or learned priors. While the early works employ simple priors such as sparsity of mass (Kim et al., 2022) and smoothness of depth (Niemeyer et al., 2022), later works found that the use of richer priors can lead to better performance. The most popular prior is the depth prior (Deng et al., 2022; Roessle et al., 2022; Uy et al., 2023; Somraj et al., 2023), while other priors based on visibility (Somraj and Soundararajan, 2023), semantic information (Jain et al., 2021) and diffusion models (Wynn and Turmukhambetov, 2023; Wu et al., 2024) are also explored. While most of the works employ the NeRF baseline leading to long training time, recent works explore the use of explicit models (Shi et al., 2024; Xiong et al., 2023; Zhu et al., 2023). However, none of these approaches analyze or regularize the distortions caused by motion in a dynamic scene.
Figure 1. Model architecture: We decompose the dynamic radiance field into a 4D scene flow or deformation field $\mathcal{F}_{f}$ that maps a 3D point $\mathbf{p}_{i}$ at time $t$ to the corresponding 3D point $\mathbf{p}_{i}^{\prime}$ at canonical time $t^{\prime}$, and a 5D radiance field $\mathcal{F}_{s}$ that models the scene at canonical time $t^{\prime}$. Both the fields are modeled using a factorized volume followed by a tiny MLP, which allows fast optimization and rendering. We note that $\mathbf{G}_{f}$ is modeled using six planes, although we show only three owing to the difficulty in visualizing four dimensions. The MLP $\mathbf{M}_{s}$ is conditioned on time and viewing direction to model time-dependent color variations such as shadows and view-dependent color variations such as specularities. The output of $\mathcal{F}_{s}$ is volume rendered to obtain the color of the pixel and the photometric loss is used to train both the fields. The explicitly modeled motion field $\mathcal{F}_{f}$ is additionally regularized using the flow priors as shown in Fig. 2.
2.5. Fast Optimization of Radiance Fields
Fast optimization of radiance fields allows wider usage in real-world applications. Voxel grid based approaches trade memory for speed by replacing the MLP in NeRF with a voxel grid (Fridovich-Keil et al., 2022; Sun et al., 2022). Prior work on fast and compact representation for learning radiance fields can be broadly classified into two categories. One category of models employs a factorized volume representation (Chen et al., 2022; Fridovich-Keil et al., 2023; Cao and Johnson, 2023) to exploit the spatial correlation of the entities to be learned, such as volume density and color. The other category of models (Kerbl et al., 2023; Müller et al., 2022; Barron et al., 2023) mainly exploit the sparsity of the scene to reduce the memory footprint. However, the scene flow for a moving object is non-zero at every time instant, but its variation is temporally correlated. Since the motion field is not sparse temporally, we believe that the sparsification based approaches may tend to learn an independent motion field for every time instant. For example, with a 4DGS model, a single 4D Gaussian may not be able to model the motion across a few time instants and can model the motion at a single time instant only, similar to Lee et al. (2023). On the other hand, factorized volumes can effectively exploit the spatio-temporal correlation of the motion field and hence we employ a 4D factorized model to learn the motion field.
3. Method
Given a set of posed video frames $I_{t}^{v}$, where $t=\left\{1,2,\ldots,N_{f}\right\}$ denotes the time instant or the frame index among $N_{f}$ frames per camera and $v=\left\{1,2,\ldots N_{c}\right\}$ denotes the camera index among $N_{c}$ stationary cameras, the goal is to synthesize the dynamic scene at a specified time instant $t=\left\{1,2,\ldots,N_{f}\right\}$ and in any novel view. We focus on the sparse multi-view setting, where the number of cameras $N_{c}$ is small. The main challenges here include the design of an explicit motion model that allows regularization with motion priors and the choice of reliable motion priors itself when training with few input views. We first describe the design of our dynamic radiance field (Sec. 3.1), and then discuss its training with reliable flow priors (Sec. 3.2).
3.1. Dynamic Radiance Field
We modify the K-Planes (Fridovich-Keil et al., 2023) implementation for dynamic radiance fields to incorporate an explicit motion model. K-Planes employs a multi-resolution grid $\mathbf{G}$ followed by a tiny multi-layer perceptron (MLP) $\mathbf{M}$ to represent a radiance field. As shown in Fig. 1, we model our dynamic radiance field by employing two factorized tensorial models, a radiance field $\mathcal{F}_{s}=\mathbf{G}_{s}\circ\mathbf{M}_{s}$ that represents the 3D scene at a canonical time instant $t^{\prime}$, and a 4D scene flow or deformation field $\mathcal{F}_{f}=\mathbf{G}_{f}\circ\mathbf{M}_{f}$ that represents the scene flow for a 3D point $\mathbf{p}$ from time $t$ to $t^{\prime}$. Mapping the 3D scene at every time instant to a canonical volume helps our model enforce temporal consistency of objects in the scene (Pumarola et al., 2021).
Here, $\mathbf{G}_{f}$ consists of six planes $\{\mathbf{S}_{xy},\mathbf{S}_{yz},\mathbf{S}_{xz},\mathbf{S}_{xt},\mathbf{S}_{yt},\mathbf{S}_{zt}\}$ at every resolution, where the first three planes model the spatial correlation and the next three model the spatio-temporal correlation of the motion field. To obtain the scene flow for a 3D point $\mathbf{p}$ from time $t$ to $t^{\prime}$, we first project the 4D point $(\mathbf{p},t)$ onto the six planes and bilinearly interpolate the feature vectors in each of the six planes. We then combine the features using the Hadamard product to obtain the final feature vector $\mathbf{h}_{i}$ as
(1) $\mathbf{h}_{i}=\mathbf{G}_{f}\left(\mathbf{p}\right)$ $=\prod_{c\in\{xy,yz,xz,xt,yt,zt\}}S_{c}\left(\mathbf{p}\right),$
which is then fed to the tiny MLP $\mathbf{M}_{f}$ that outputs the scene flow. Modeling the scene flow field $\mathcal{F}_{f}$ using the hex-plane representation and the tiny MLP makes our motion model fast and compact. For more details on the hex-plane representation, we refer the readers to K-Planes (Fridovich-Keil et al., 2023). We model the scene grid $\mathbf{G}_{s}$ using a similar factorization with three spatial planes. While it is beneficial to model the motion field using a factorized volume, the canonical scene can be learned using any explicit representation such as 3DGS (Kerbl et al., 2023). We employ factorized volumes to model both scene and motion using a unified framework.
We train our model, similar to K-Planes, by rendering randomly sampled pixels $\mathbf{q}_{t}^{v}$ in view $v$ at time $t$ and using the ground truth color as supervision. To render a pixel $\mathbf{q}_{t}^{v}$, we sample $N_{p}$ points $\left\{\mathbf{p}_{i}\right\}_{i=1}^{N_{p}}$ at depths $\left\{z_{i}\right\}_{i=1}^{N_{p}}$ along the corresponding ray. For every 3D point $\mathbf{p}_{i}$, we first obtain the corresponding 3D point $\mathbf{p}_{i}^{\prime}$ at canonical time $t^{\prime}$ by computing the scene flow from time $t$ to $t^{\prime}$ using $\mathcal{F}_{f}$ as
(2) $\mathbf{p}_{i}^{\prime}$ $=\mathcal{F}_{f}\left(\mathbf{p}_{i},t\right)+\mathbf{p}_{i}.$
We then query $\mathbf{G}_{s}$ at $\mathbf{p}_{i}^{\prime}$ to obtain the volume density $\sigma_{i}$ and a latent feature $\mathbf{h}_{i}^{\prime}$ corresponding to $\mathbf{p}_{i}$ as
(3) $\sigma_{i},\mathbf{h}_{i}^{\prime}=\mathbf{G}_{s}\left(\mathbf{p}_{i}^{\prime}\right).$
A tiny MLP $\mathbf{M}_{s}$ maps $\mathbf{h}_{i}^{\prime}$, encoded viewing direction $\mathbf{v}$ and encoded time $t$ to the color $\mathbf{c}_{i}$ of $\mathbf{p}_{i}$ as
(4) $\mathbf{c}_{i}=\mathbf{M}_{s}\left(\mathbf{h}_{i}^{\prime},\gamma\left(\mathbf{v}\right),\gamma\left(t\right)\right),$
where $\gamma$ denotes the encoding (Müller et al., 2022) of the viewing direction and the time instant. Conditioning $\mathbf{M}_{s}$ additionally on time allows our model to capture illumination changes due to object motion. $\mathbf{c}_{i}$ are then volume rendered to obtain the color $\mathbf{c}$ of $\mathbf{q}$ as
(5) $w_{i}$ $=\left(\Pi_{j=1}^{i-1}\exp\left(-\delta_{j}\sigma_{j}\right)\right)\left(1-\exp\left(-\delta_{i}\sigma_{i}\right)\right),$
(6) $\mathbf{c}$ $=\sum_{i=1}^{N_{p}}w_{i}\mathbf{c}_{i},\quad z=\sum_{i=1}^{N_{p}}w_{i}z_{i},$
where $\delta_{i}=z_{i}-z_{i-1}$ and $z$ gives the expected depth of $\mathbf{q}_{t}^{v}$. Both $\mathcal{F}_{f}$ and $\mathcal{F}_{s}$ are optimized through the photometric loss, ${\mathcal{L}_{\text{ph}}}=\|\mathbf{c}-\hat{\mathbf{c}}\|^{2}$, where $\hat{\mathbf{c}}$ is the ground truth color. Decomposing the scene into a canonical scene field and a flow field allows us to regularize the flow field using flow priors when only a sparse set of viewpoints are available.
Figure 2. Flow regularization: Since the motion field $\mathcal{F}_{f}$ gives only the unidirectional flow from time $t$ to $t^{\prime}$, we impose the flow prior by minimizing the distance between the 3D points in the canonical volume corresponding to the matched pixels $(\mathbf{q}_{t}^{v},\mathbf{q}_{s}^{u})$ in the input frames $(I_{t}^{v},I_{s}^{u})$.
3.2. Training with Flow Priors
(a) Dense flow across cameras estimated by RAFT.
(b) Sparse flow across cameras estimated by SIFT.
(c) Dense flow within cameras estimated by RAFT.
Figure 3. Visualization of different flow priors: We show the matched pixels as provided by different flow priors. The pixels in the first view are randomly picked from those for which sparse flow is available and the same pixels are used for dense flow. Note that the second view in the first two examples has more blur as compared to the first view. (a) We show that the dense flow priors across cameras obtained using deep optical flow networks such as RAFT (Teed and Deng, 2020) are prone to erroneous matches, due to variations in camera parameters and lighting. We observed similar trends with other deep optical flow networks as well such as AR-Flow (Liu et al., 2020). (b) Matching pixels across cameras using robust SIFT features provides reliable matches, albeit sparse. (c) Within individual cameras, the dense correspondences provided by deep optical flow networks are more reliable owing to smaller variations in lighting.
N3DV InterDigital
Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$
2 views HexPlane 13.81 0.41 0.58 2.03 13.78 0.24 0.55 1.51
K-Planes 17.95 0.62 0.43 0.59 16.51 0.51 0.40 0.35
DeRF 17.98 0.61 0.44 0.67 17.04 0.56 0.37 0.35
RF-DeRF 20.35 0.71 0.33 0.41 19.08 0.66 0.31 0.23
3 views HexPlane 15.53 0.49 0.50 1.97 13.85 0.27 0.54 1.43
K-Planes 23.65 0.83 0.25 0.34 18.75 0.66 0.30 0.22
DeRF 22.61 0.81 0.27 0.39 18.72 0.68 0.30 0.22
RF-DeRF 25.07 0.88 0.21 0.16 20.41 0.74 0.25 0.12
4 views HexPlane 15.50 0.47 0.54 1.95 15.18 0.38 0.49 1.24
K-Planes 24.80 0.86 0.20 0.27 23.85 0.86 0.16 0.11
DeRF 23.76 0.83 0.24 0.36 23.51 0.85 0.17 0.13
RF-DeRF 25.86 0.89 0.19 0.17 23.69 0.86 0.17 0.10
Table 2. Quantitative results: We compare our model with K-Planes on N3DV (Li et al., 2022) and InterDigital (Sabater et al., 2017) datasets with two, three, and four input views. We also show the performance of our base DeRF model for reference. We report the PSNR, SSIM, and LPIPS scores for the rendered images and the depth MAE for the rendered depth maps. The best scores in each category are shown in bold.
A dense flow prior across cameras and across time instants is an ideal prior for optimal regularization of our motion field. However, different cameras in a multi-view setting could suffer from variations in parameters such as intrinsics, color balance, exposure and focus making it challenging to determine correspondences across cameras based on luminance constancy. As a result, the dense flow estimates obtained using deep optical flow networks are unreliable across cameras as seen in Fig. 3(a). On the other hand, matching keypoints using robust SIFT (Lowe, 2004) descriptors are more reliable across cameras as seen in Fig. 3(b). Specifically, for a pair of frames across time instants $t$ and $s$ and cameras $v$ and $u$, we extract the SIFT keypoints in individual images using Colmap (Schonberger and Frahm, 2016) and match the keypoints using SIFT descriptors. The difference between the locations of the matched keypoints gives the sparse flow prior $\mathcal{P}_{\text{sf}}$. While Colmap provides a reasonable number of matches, newer models such as R2D2 (Revaud et al., 2019) could be employed to improve the richness of the sparse flow prior.
Employing flow priors from any time $t$ to canonical time $t^{\prime}$ for large $t-t^{\prime}$ may be less efficient since large regions of the scene may not be visible. Thus, it is desirable to utilize flow priors across a short duration of time. However, the motion model $\mathcal{F}_{f}$ provides the scene flow from $t$ to $t^{\prime}$ only and not the flow between any two arbitrary time instants. We resolve this by using the flow prior to encourage $\mathcal{F}_{f}$ to map a pair of matched points at different time instants to the same object location in 3D. Specifically, we obtain the matching pixels using the flow priors and constrain the 3D points corresponding to the matching pixels to map to the same 3D point in the canonical scene field as shown in Fig. 2.
Mathematically, let the flow corresponding to a pixel $\mathbf{q}_{t}^{v}$ to time $s$ and view $u$ be given by $\mathbf{f}_{t\rightarrow s}^{v\rightarrow u}$. Then the matching pixel at time $s$ and view $u$ is given by $\mathbf{q}_{s}^{u}=\mathbf{q}_{t}^{v}+\mathbf{f}_{t\rightarrow s}^{v\rightarrow u}$. Let the 3D points sampled along the rays corresponding to $\mathbf{q}_{t}^{v}$ and $\mathbf{q}_{s}^{u}$ be given by $\mathbf{p}_{i}(t,v)$ and $\mathbf{p}_{i}(s,u)$ respectively. Then, we impose the sparse flow constraint as
(7) ${\mathcal{L}_{\text{sf}}}$ $=\left\|\sum_{i=0}^{N_{p}-1}w_{i}(t,v)\mathbf{p}_{i}^{\prime}(t,v)-\sum_{i=0}^{N_{p}-1}w_{i}(s,u)\mathbf{p}_{i}^{\prime}(s,u)\right\|^{2},$
where $\mathbf{p}^{\prime}(t,v)$ and $\mathbf{p}^{\prime}(s,u)$ are computed from $\mathbf{p}(t,v)$ and $\mathbf{p}(s,u)$ using $\mathcal{F}_{f}$ as in Eq. 2, and $w_{i}(t,v)$ and $w_{i}(s,u)$ are computed using Eq. 5. The two terms in Eq. 7 represent $\mathbf{q}_{t}^{v}$ and $\mathbf{q}_{s}^{u}$ in the canonical volume respectively. Thus, ${\mathcal{L}_{\text{sf}}}$ guides $\mathcal{F}_{s}$ on which two points in two different time instants belong to the same object.
In Eq. 7, we first find the canonical field points $\mathbf{p}_{i}^{\prime}$ using $\mathcal{F}_{f}$ and then average the location of the 3D points using weights $w_{i}$ and not the other way for the following reason. The former approach regularizes the motion model $\mathcal{F}_{f}$ for every $\mathbf{p}_{i}$ giving a rich supervision to $\mathcal{F}_{s}$, whereas the latter approach regularizes $\mathcal{F}_{f}$ only for the expected 3D point $\mathbf{p}=\sum w_{i}\mathbf{p}_{i}$. Further, since we do not impose stop-gradient on $w_{i}$ in ${\mathcal{L}_{\text{sf}}}$, the flow priors also help remove incorrect masses such as floaters.
While the sparse flow prior $\mathcal{P}_{\text{sf}}$ is rich on account of matching pixels across cameras and time instants, the prior is available at only sparse keypoints. Thus, we complement $\mathcal{P}_{\text{sf}}$ with a dense flow prior $\mathcal{P}_{\text{df}}$ within individual cameras. For a pair of frames from the same camera at two closer time instants, there is little variation in the camera intrinsics and pipeline parameters. Hence, the dense flow priors obtained using a deep optical flow network are more reliable as seen in Fig. 3(c). We employ the popular optical flow estimation model RAFT (Teed and Deng, 2020) to obtain the dense flow priors and constrain our model with a dense flow loss ${\mathcal{L}_{\text{df}}}$ similar to Eq. 7. We find that the dense flow priors typically help reduce artifacts in static regions.
We train our model by minimizing the combination of photometric loss ${\mathcal{L}_{\text{ph}}}$, sparse flow loss ${\mathcal{L}_{\text{sf}}}$ and dense flow loss ${\mathcal{L}_{\text{df}}}$ as
(8) ${\mathcal{L}}$ $={\mathcal{L}_{\text{ph}}}+\lambda_{\text{sf}}{\mathcal{L}_{\text{sf}}}+\lambda_{\text{df}}{\mathcal{L}_{\text{df}}},$
where $\lambda_{\text{sf}}$ and $\lambda_{\text{df}}$ are hyper-parameters.
Figure 4. Qualitative examples on N3DV dataset with two input views: We can observe that K-Planes finds it hard to learn the moving person leading to significant distortions. Our DeRF model (without any priors) corrects a few errors by virtue of the common canonical volume. Imposing our priors leads to much better reconstruction.
4. Experiments
4.1. Evaluation Setup
Datasets: We evaluate our model on two popular multi-view dynamic scene datasets, namely N3DV (Li et al., 2022) and InterDigital (Sabater et al., 2017) with two, three, and four input views. Following prior work (Fridovich-Keil et al., 2023), we downsample the videos spatially by a factor of two for all the experiments. We use the video at the center of the camera rig for testing and uniformly sample train videos from the remaining videos. N3DV dataset contains six real-world scenes with 17–21 static cameras per scene and 300 frames per viewpoint. The videos have a spatial resolution of $1352\times 1014$ and a frame rate of 30fps. InterDigital dataset contains multiple real-world scenes with 16 static cameras and varying number of frames per scene. We undistort the video frames using the radial distortion parameters provided with the dataset and use the undistorted videos for our experiments. We select five scenes that contain at least 300 frames and choose the first 300 frames. The videos have a spatial resolution of $1024\times 544$ and a frame rate of 30fps spanning 10 seconds in all our experiments.
Evaluation measures: We evaluate the rendered frames of all the methods using PSNR, SSIM (Wang et al., 2003) and LPIPS (Zhang et al., 2018). We also evaluate the models on their ability to reconstruct the 3D scene by computing MAE on the rendered depth maps. Due to the unavailability of true depth maps on both datasets, we use the depth provided by K-Planes trained with dense input views as pseudo ground truth depth.
$\mathcal{P}_{\text{sf}}$ $\mathcal{P}_{\text{df}}$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$
22.61 0.81 0.27 0.39
24.79 0.87 0.22 0.20
24.10 0.85 0.23 0.31
25.07 0.88 0.21 0.16
Table 3. Quantitative results of ablations on N3DV dataset with three input views: We evaluate the significance of the two flow priors we employ, namely the sparse flow prior ($\mathcal{P}_{\text{sf}}$) and the dense flow prior ($\mathcal{P}_{\text{df}}$) by disabling one component at a time. We also show the performance of the DeRF model without any priors for reference.
4.2. Comparisons and Implementation Details
We mainly compare the performance of our model against K-Planes and HexPlane. We use the official code released by the authors of K-Planes and modify it to implement our model. We refer to our base model without any priors as DeRF model. To test the superiority of our priors, we also compare our model against K-Planes with sparse depth priors and DeRF with dense flow priors across cameras. Since we use two factorized models instead of one used in K-Planes, we reduce the feature dimension in both of our models by half to keep the total number of parameters comparable to K-Planes. We train all the models for 30k iterations on a single NVIDIA RTX 2080 GPU. We randomly pick $s\in\{t-10,t+10\}$ to impose the flow prior losses and set $\lambda_{\text{sf}}=1$ and $\lambda_{\text{df}}=1$. We use the same values as suggested by K-Planes for all the remaining hyperparameters.
N3DV InterDigital
Model LPIPS $\downarrow$ Depth MAE $\downarrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$
w/o any priors 0.27 0.39 0.30 0.22
w/ cross-camera dense flow 0.30 0.49 0.36 0.29
w/ our flow priors 0.21 0.16 0.25 0.12
Table 4. We test the performance of our DeRF model with dense flow priors across cameras and our reliable flow priors. The performance of the DeRF model (without any priors) is also shown for reference.
4.3. Results
We show the quantitative performance of our model, K-Planes and HexPlane in Tab. 2, where we observe that our RF-DeRF model outperforms both K-Planes and HexPlane across all the settings on both the datasets. We observe that the performance of all the models is relatively higher on the N3DV dataset as compared to the InterDigital dataset. This is perhaps due to the InterDigital dataset having larger motion and highly textured regions. We also note that the performance of HexPlane is substantially lower than both K-planes and our model. This could be a result of optimizing HexPlane initially with a low-resolution grid, which causes the model to overfit the input-views and not utilize the high-resolution grid in the later stages of training. K-Planes does not suffer from this drawback by using a multi-resolution grid throughout the optimization. From Figs. 6, 7 and 9, we observe that our model is able to correct errors both in moving objects and in static regions over K-Planes. In particular, we observe that K-Planes suffer from blur, deformation, and disappearance of moving objects while rendering novel views, which are mitigated by our model. The improvements are more starkly visible in the supplementary videos. We compare the rendered depth from both the models in Figs. 11 and 11 and observe that our model is able to reconstruct the depth more accurately than K-Planes. Tab. 2 also shows that our DeRF model (without any priors) performs better than K-Planes in the scenes with larger motion, such as the scenes in the InterDigital dataset, perhaps due to the temporal consistency enforced by the canonical volume. This is also visible in the example shown in Fig. 4. In our experiments, RF-DeRF roughly took 1.5 hours and 5GB GPU memory to train on a single scene, and 6 seconds to render a single frame. The size of the saved model parameters is approximately 280MB. Please refer to Fridovich-Keil et al. (2023) for the training time for various models including K-Planes and DyNeRF.
Figure 5. Qualitative examples for depth priors vs flow priors: We observe that sparse depth prior is not very effective in improving the reconstruction quality. However, our sparse flow prior is highly effective in mitigating the distortions, while the use of sparse and dense flow priors in our final model leads to the best reconstruction quality.
Ablations: We analyze the significance of each component of our model by disabling one component at a time in Tab. 3. We observe that the sparse flow prior is the most significant component of our model, giving a large boost in performance when included and a large drop in performance when excluded. The effect of sparse flow prior is more pronounced in terms of depth MAE, which measures the accuracy of the reconstructed 3D scene. This demonstrates the importance of reliable motion priors across cameras, even if sparse. Fig. 8 shows that sparse flow prior is more helpful in mitigating motion related artifacts such as motion blur, while within-camera dense flow prior mainly helps reduce distortions in the static regions. The combination of the two priors leads to the best performance by mitigating both artifacts.
N3DV InterDigital
Model SSIM $\uparrow$ Depth MAE $\downarrow$ SSIM $\uparrow$ Depth MAE $\downarrow$
K-Planes 0.83 0.34 0.66 0.22
K-Planes + $\mathcal{P}_{\text{sd}}$ 0.82 0.26 0.71 0.20
DeRF 0.81 0.39 0.68 0.22
DeRF + $\mathcal{P}_{\text{sd}}$ 0.80 0.34 0.70 0.22
DeRF + $\mathcal{P}_{\text{sf}}$ 0.87 0.20 0.73 0.13
RF-DeRF 0.88 0.16 0.74 0.12
Table 5. We analyze the impact of our sparse flow prior that regularizes both the scene geometry and the learned motion against the sparse depth prior (Deng et al., 2022) that regularizes only the scene geometry. The models are trained with three input views on both the datasets.
Dense cross-camera flow priors vs our priors: To validate our hypothesis that a combination of cross-camera sparse flow and within-camera dense flow priors is more effective than dense flow priors across cameras, we impose cross-camera dense flow priors on our base DeRF model and evaluate the performance on both datasets. From Tab. 4 and Fig. 13, we observe that imposing noisy cross-camera dense flow priors leads to a large drop in performance, whereas our priors help improve the performance significantly.
Depth vs flow priors: We now analyze the effect of flow priors against that of depth priors in dynamic view synthesis with sparse input views. The depth priors guide the model on the exact locations of the objects in the scene, but do not account for the correlation of object locations across time instants. On the other hand, the flow priors account for such inter-frame correlations as well as encompass the information provided by the depth priors in static regions. To experimentally validate our hypothesis, we compare the performance of K-Planes and our DeRF models when regularized with sparse depth priors $\mathcal{P}_{\text{sd}}$ obtained through Colmap sparse reconstruction (Deng et al., 2022). We obtain $\mathcal{P}_{\text{sd}}$ for every frame and use them to constrain the K-Planes model as ${\mathcal{L}_{\text{sd}}}=\|z-\hat{z}\|^{2}$, where $z$ is the depth obtained in Eq. 6 and $\hat{z}$ is the depth prior. From Tab. 5, we observe that sparse flow priors lead to a significant improvement in performance as compared to that of sparse depth priors. The difference in performance can be more clearly observed in Fig. 5 and the supplementary videos. This shows the utility of motion priors in dynamic view synthesis with sparse input views.
Performance with increasing viewpoints: To understand the impact of our priors as the number of input views increases, we compare the performance of RF-DeRF against K-Planes with different number of cameras on a single scene from N3DV dataset in Fig. 13. Our model significantly improves the performance over K-Planes in sparse viewpoint scenarios, while matching the performance of K-Planes with dense input views. Thus, we conclude that our priors are particularly helpful in improving the reconstruction of under-observed regions in the scene.
N3DV InterDigital
Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
K-Planes 30.55 0.96 0.12 29.00 0.96 0.07
DeRF 29.95 0.95 0.12 28.14 0.95 0.09
Table 6. We test the performance of our DeRF model against K-Planes with dense input views. Please see the videos in the supplementary for visual comparisons.
Performance with dense input views: We now test the ability of our model in reconstructing the dynamic scene when provided with dense multi-view inputs. Since motion priors may not be needed when dense views are available, we analyze the performance of our base DeRF model (without any priors). From Tab. 6, we observe that our DeRF model is competitive with K-Planes. Thus, one could employ our model in both dense and sparse input scenarios by additionally employing reliable flow priors in the latter case.
4.4. Limitations and Future Work
Our approach of mapping the scene at every time instant to the canonical volume implies that only the objects present in the canonical volume can be rendered. In other words, our model may not handle objects entering or leaving the scene, which could happen over longer durations. This could be resolved by learning multiple models, each trained on self-contained shorter duration videos. We employ Colmap to generate sparse flow priors in our framework, which is a time-consuming process. In our experiments, we found that the generation of sparse flow priors takes about 45 minutes per scene. There is a need to explore faster approaches to determine reliable sparse correspondences across cameras. Further, our approach requires the cameras to be calibrated and time-synchronized apriori. It would be interesting to optimize the scene, camera parameters, and time-synchronization jointly in a single framework.
5. Conclusion
We consider the setting of fast dynamic view synthesis when only a few videos of the scene as observed from different static cameras are available. By exploiting the spatio-temporal correlation of the motion field, we design an explicit motion model using factorized representations that is compact, fast, and allows effective regularization with the flow priors. We observe that the naive use of cross-camera dense flow priors has a negative effect on the performance, while a careful imposition of reliable cross-camera sparse and within-camera dense flow priors provides a significant boost in performance. Further, we also show that the motion priors are more general and encompass the depth priors in dynamic view synthesis with sparse input views. We demonstrate the effectiveness of our approach on two popular datasets and show that our approach outperforms the state-of-the-art fast and compact dynamic radiance fields by a large margin when only a few viewpoints are available.
Acknowledgements.
This work was supported in part by a grant from Kotak IISc AI/ML centre (KIAC). The first author was supported partially by the Prime Minister’s Research Fellowship awarded by the Ministry of Education, Government of India.
Supplement
The contents of this supplement include
  1. A.
    Video examples on N3DV and InterDigital datasets.
  2. B.
    Extensive quantitative evaluation reports.
Appendix A Video Comparisons
We compare our RF-DeRF model with K-Planes (Fridovich-Keil et al., 2023), ablated models, and various other baselines we design. We divide the video comparisons into three sets. In the first set, we compare our model against the vanilla K-Planes and the different baselines we design by incorporating sparse-depth and cross-camera dense flow priors. In the second set, we show the effect of removing each of our priors. Finally, we compare our base DeRF (without any priors) against K-Planes when dense input views are available. The videos are available on our project website https://nagabhushansn95.github.io/publications/2024/RF-DeRF.html.
Appendix B Performance on Individual Scenes
For the benefit of follow-up work, where researchers may want to analyze the performance of different models or compare the models on individual scenes, we provide the performance of various models on individual scenes in Tabs. 8, 7, 9, 11, 10 and 12.
Table 7. Per-scene performance of various models with three input views on N3DV dataset. The four rows show PSNR, SSIM, LPIPS Alex, and Depth MAE scores, respectively.
Model \ Scene Name Coffee Martini Cook Spinach Cut Roasted Beef Flame Salmon Flame Steak Sear Steak Average
HexPlane 13.70 0.3892 0.5814 1.985 15.72 0.5109 0.4827 2.0593 18.40 0.6247 0.3937 1.6426 11.03 0.3109 0.6558 2.3403 16.94 0.5523 0.451 1.9577 17.42 0.5705 0.4296 1.8327 15.53 0.4931 0.499 1.9696
K-Planes 20.11 0.7590 0.3215 0.4501 24.24 0.8743 0.2059 0.3275 26.55 0.9128 0.1767 0.1570 16.92 0.5975 0.4717 0.7002 26.77 0.9178 0.1541 0.2178 27.29 0.9216 0.1520 0.1718 23.65 0.8305 0.2470 0.3374
K-Planes + Sparse Depth 20.66 0.7830 0.2672 0.2910 22.70 0.8557 0.1916 0.2468 26.66 0.9090 0.1541 0.1058 17.19 0.6167 0.4149 0.5459 24.52 0.8767 0.1700 0.2011 25.02 0.8971 0.1486 0.1640 22.79 0.8230 0.2244 0.2591
DeRF 19.82 0.7531 0.3282 0.4619 23.15 0.8361 0.2338 0.3320 25.19 0.8946 0.2018 0.2080 16.93 0.5852 0.4579 0.6320 25.33 0.8819 0.1917 0.3483 25.25 0.8805 0.1909 0.3340 22.61 0.8052 0.2674 0.3860
DeRF + Sparse Depth 19.81 0.7741 0.3088 0.3694 22.64 0.8316 0.2503 0.3332 25.29 0.8813 0.2105 0.1473 16.47 0.5572 0.4732 0.6096 24.59 0.8676 0.2199 0.3325 24.85 0.8783 0.1983 0.2484 22.27 0.7983 0.2768 0.3401
DeRF + Cross-Cam Dense Flow 19.52 0.6990 0.4035 0.6330 23.95 0.8439 0.2443 0.4498 24.94 0.8718 0.2402 0.3664 17.73 0.6043 0.4929 0.7207 25.24 0.8776 0.2055 0.3959 25.24 0.8832 0.1990 0.3937 22.77 0.7966 0.2976 0.4932
RF-DeRF w/o Sparse Flow 20.58 0.8016 0.2923 0.3089 25.61 0.9052 0.1808 0.2936 26.79 0.9174 0.1835 0.2293 17.93 0.6680 0.3944 0.5329 27.07 0.9233 0.1533 0.2365 26.62 0.9099 0.1687 0.2569 24.10 0.8542 0.2288 0.3097
RF-DeRF w/o Dense Flow 21.10 0.8106 0.2863 0.2648 26.03 0.9003 0.1875 0.1791 27.55 0.9225 0.1760 0.1254 19.56 0.7452 0.3622 0.2866 26.85 0.9153 0.1671 0.2060 27.66 0.9280 0.1518 0.1457 24.79 0.8703 0.2218 0.2013
RF-DeRF 21.63 0.8233 0.2700 0.2095 26.29 0.9043 0.1850 0.1601 26.67 0.9135 0.1925 0.1620 20.77 0.7776 0.3194 0.1877 27.95 0.9330 0.1541 0.1320 27.11 0.9192 0.1669 0.1345 25.07 0.8785 0.2146 0.1643
Table 8. Per-scene performance of various models with two input views on N3DV dataset. The four rows show PSNR, SSIM, LPIPS Alex, and Depth MAE scores, respectively.
Model \ Scene Name Coffee Martini Cook Spinach Cut Roasted Beef Flame Salmon Flame Steak Sear Steak Average
HexPlane 13.10 0.3984 0.5565 2.0923 14.14 0.4284 0.5848 2.0748 15.16 0.4507 0.5165 1.6882 10.56 0.2518 0.6963 2.3864 15.04 0.4735 0.5673 2.0209 14.84 0.4707 0.5504 1.9361 13.81 0.4123 0.5786 2.0331
K-Planes 16.20 0.5295 0.4553 0.7790 17.56 0.5966 0.4775 0.5976 22.32 0.8413 0.2253 0.2024 13.36 0.3590 0.6570 0.9757 18.88 0.6899 0.3967 0.4730 19.36 0.7044 0.3683 0.4822 17.95 0.6201 0.4300 0.5850
K-Planes + Sparse Depth 15.87 0.5396 0.4477 0.7115 17.34 0.6239 0.4264 0.4893 20.94 0.8078 0.2547 0.2193 12.65 0.3563 0.6167 0.7521 17.65 0.6310 0.4151 0.5576 17.66 0.6658 0.3807 0.4673 17.02 0.6041 0.4235 0.5328
DeRF 15.84 0.5141 0.4706 0.8747 17.68 0.6070 0.4591 0.6107 22.78 0.8156 0.2650 0.3514 12.57 0.3481 0.6424 1.0873 19.02 0.6640 0.4100 0.5580 19.99 0.7030 0.3849 0.5492 17.98 0.6086 0.4386 0.6719
RF-DeRF 17.30 0.6005 0.4165 0.5743 21.74 0.7707 0.3046 0.3404 24.16 0.8731 0.1861 0.1971 13.67 0.4097 0.6103 0.7844 22.70 0.8079 0.2535 0.3119 22.51 0.8190 0.2381 0.2759 20.35 0.7135 0.3349 0.4140
Table 9. Per-scene performance of various models with four input views on N3DV dataset. The four rows show PSNR, SSIM, LPIPS Alex, and Depth MAE scores, respectively.
Model \ Scene Name Coffee Martini Cook Spinach Cut Roasted Beef Flame Salmon Flame Steak Sear Steak Average
HexPlane 13.99 0.3846 0.5682 1.9928 14.52 0.4299 0.5939 2.0695 21.92 0.8078 0.2289 1.3548 11.21 0.316 0.6671 2.3662 15.76 0.4482 0.5837 2.0197 15.62 0.4429 0.5701 1.9188 15.50 0.4716 0.5353 1.9536
K-Planes 22.89 0.8603 0.1971 0.3168 24.71 0.8702 0.1911 0.2588 28.64 0.9418 0.1068 0.1506 18.91 0.6985 0.3657 0.5428 26.90 0.9040 0.1632 0.1578 26.73 0.9051 0.1510 0.1730 24.80 0.8633 0.1958 0.2666
K-Planes + Sparse Depth 20.26 0.7938 0.2646 0.2291 21.08 0.7696 0.2957 0.3044 25.69 0.9029 0.1607 0.1493 18.14 0.6907 0.3608 0.2761 21.28 0.7564 0.2959 0.3128 22.47 0.8052 0.2454 0.2662 21.49 0.7864 0.2705 0.2563
DeRF 22.02 0.8468 0.2290 0.3545 23.36 0.8167 0.2603 0.3749 27.91 0.9295 0.1306 0.1590 17.70 0.6273 0.4303 0.6675 25.72 0.8679 0.2069 0.3359 25.86 0.8857 0.1794 0.2631 23.76 0.8290 0.2394 0.3591
RF-DeRF 23.43 0.8798 0.2089 0.1536 25.76 0.8699 0.2101 0.2093 28.32 0.9310 0.1364 0.1523 22.85 0.8629 0.2329 0.1524 27.10 0.8979 0.1722 0.2077 27.68 0.9070 0.1621 0.1685 25.86 0.8914 0.1871 0.1740
Table 10. Per-scene performance of various models with three input views on InterDigital dataset. The four rows show PSNR, SSIM, LPIPS Alex, and Depth MAE scores, respectively.
Model \ Scene Name Birthday Painter Remy Theater Train Average
HexPlane 12.54 0.2115 0.5298 1.2777 17.42 0.38 0.433 1.4585 13.34 0.4295 0.4888 1.1883 14.75 0.3153 0.5735 1.3894 11.24 0.0001 0.6867 1.8348 13.86 0.2673 0.5424 1.4297
K-Planes 19.58 0.8639 0.1388 0.0886 26.42 0.9369 0.0892 0.0700 15.18 0.4302 0.5086 0.2936 17.01 0.5339 0.3949 0.2530 15.56 0.5591 0.3918 0.3846 18.75 0.6648 0.3047 0.2179
K-Planes + Sparse Depth 19.49 0.8476 0.1525 0.1194 26.26 0.9305 0.0977 0.1708 15.59 0.4929 0.4580 0.2091 17.59 0.5744 0.3737 0.3036 17.84 0.6870 0.2792 0.2189 19.36 0.7065 0.2722 0.2044
DeRF 19.55 0.8636 0.1533 0.0987 25.93 0.9272 0.1037 0.1344 16.39 0.5377 0.4413 0.2121 16.69 0.5248 0.4144 0.2608 15.02 0.5443 0.3990 0.3938 18.72 0.6795 0.3024 0.2200
DeRF + Sparse Depth 19.25 0.8441 0.1615 0.1149 25.98 0.9249 0.1107 0.2323 15.34 0.5037 0.4566 0.2335 16.02 0.5338 0.4081 0.2832 18.03 0.6691 0.2908 0.2155 18.92 0.6951 0.2855 0.2159
DeRF + Cross-Cam Dense Flow 18.25 0.7215 0.3034 0.2503 24.56 0.8613 0.1802 0.3576 16.20 0.5050 0.4735 0.3198 16.10 0.4545 0.4761 0.2397 17.13 0.6095 0.3677 0.2720 18.45 0.6304 0.3602 0.2879
RF-DeRF w/o Sparse Flow 19.26 0.8339 0.1946 0.1624 25.91 0.9099 0.1245 0.3070 15.73 0.4918 0.4974 0.2781 17.23 0.5346 0.4108 0.2833 17.27 0.6556 0.3138 0.2126 19.08 0.6851 0.3082 0.2487
RF-DeRF w/o Dense Flow 19.51 0.8723 0.1412 0.0666 26.51 0.9397 0.0917 0.0701 17.69 0.5594 0.4154 0.2036 18.45 0.5967 0.3436 0.2112 18.14 0.6901 0.2874 0.0775 20.06 0.7316 0.2559 0.1258
RF-DeRF 19.91 0.8680 0.1449 0.0718 26.55 0.9306 0.1038 0.0953 17.51 0.5749 0.4036 0.1929 18.84 0.6092 0.3471 0.1952 19.24 0.7265 0.2546 0.0639 20.41 0.7418 0.2508 0.1238
Table 11. Per-scene performance of various models with two input views on InterDigital dataset. The four rows show PSNR, SSIM, LPIPS Alex, and Depth MAE scores, respectively.
Model \ Scene Name Birthday Painter Remy Theater Train Average
HexPlane 11.94 0.1623 0.578 1.3664 16.58 0.2151 0.4826 1.6554 13.75 0.4065 0.4829 1.2256 14.32 0.2497 0.6028 1.4564 12.31 0.1499 0.6172 1.8696 13.78 0.2367 0.5527 1.5147
K-Planes 16.71 0.6798 0.2726 0.2105 20.23 0.6687 0.2905 0.3262 14.18 0.4016 0.5047 0.3790 16.24 0.4168 0.4573 0.3061 15.20 0.3991 0.4852 0.5471 16.51 0.5132 0.4021 0.3538
DeRF 17.76 0.7572 0.2286 0.1782 21.27 0.7271 0.2487 0.3787 15.00 0.4778 0.4659 0.3275 16.12 0.4332 0.4600 0.3477 15.06 0.3958 0.4590 0.5409 17.04 0.5582 0.3725 0.3546
RF-DeRF 19.08 0.7913 0.2239 0.1688 24.38 0.8360 0.1855 0.2542 15.62 0.4952 0.4728 0.3404 17.68 0.4905 0.4050 0.2870 18.64 0.6914 0.2837 0.1124 19.08 0.6609 0.3142 0.2326
Table 12. Per-scene performance of various models with four input views on InterDigital dataset. The four rows show PSNR, SSIM, LPIPS Alex, and Depth MAE scores, respectively.
Model \ Scene Name Birthday Painter Remy Theater Train Average
K-Planes 23.26 0.9294 0.0895 0.0553 30.33 0.9708 0.0642 0.0547 21.04 0.7512 0.2747 0.1549 23.72 0.8288 0.1897 0.1475 20.93 0.8163 0.1674 0.1448 23.85 0.8593 0.1571 0.1114
DeRF 23.31 0.9243 0.1013 0.0697 29.60 0.9662 0.0725 0.0965 21.46 0.7649 0.2552 0.1431 22.45 0.7598 0.2444 0.1980 20.72 0.8114 0.1879 0.1385 23.51 0.8453 0.1723 0.1292
RF-DeRF 23.02 0.9228 0.0992 0.0519 29.22 0.9671 0.0688 0.0706 21.74 0.7571 0.2606 0.1398 22.82 0.7748 0.2484 0.1776 21.65 0.8546 0.1541 0.0632 23.69 0.8553 0.1662 0.1006